OpenStudy (michaelbp):

If (f + g)(x) = 3x2 + 2x – 1 and g(x) = 2x – 2, what is f(x)?

OpenStudy (knov):

can you post the equation using Equation option ?

OpenStudy (michaelbp):

\[(f+g)(x)=3x ^{2}+2x-1 g(x)=2x-2\]

OpenStudy (michaelbp):

space inbetween the -1 and g

zepdrix (zepdrix):

\[\large\rm \color{orangered}{(f+g)(x)=3x^2+2x-1},\qquad\qquad \color{royalblue}{g(x)=2x-2}\] \[\large\rm (f+g)(x)=f(x)+g(x)\]If you subtract g(x) from both sides,\[\large\rm (f+g)(x)-g(x)=f(x)\]

OpenStudy (knov):

am i the only one who is seeing 'math processing error' error ?

zepdrix (zepdrix):

\[\large\rm \color{orangered}{(f+g)(x)}-\color{royalblue}{g(x)}=f(x)\]\[\large\rm \color{orangered}{3x^2+2x-1}-\color{royalblue}{2x-2}=f(x)\]

OpenStudy (michaelbp):

\[3x ^{2}+4x-3\]

zepdrix (zepdrix):

Oh I didn't put parenthesis around the g(x) my mistake :)\[\large\rm \color{orangered}{3x^2+2x-1}-\color{royalblue}{(2x-2)}=f(x)\]The subtraction needs to be distributed.

zepdrix (zepdrix):

Hmm I don't think you have 4 x's, you have 2x and -2x.

OpenStudy (michaelbp):

my mistake i dont really know how to do algebra that well

zepdrix (zepdrix):

Distributing the negative gives us,\[\large\rm 3x^2+2x-1-2x+2\]Combine like-terms. 2x and -2x -1 and 2

OpenStudy (michaelbp):

\[3x ^{2}+1\]

zepdrix (zepdrix):

f(x)=3x^2+1 yay good job \c:/

OpenStudy (michaelbp):

thanks